# Thermodynamic Descriptors from Molecular Dynamics as Machine Learning Features for Extrapolable Property Prediction

### 12 Mar 2026 v1

@@ Abstract
Machine learning (ML) models which rely on molecular structure excel at predicting properties for well-represented organic compounds, however their limited ability to extrapolate to chemotypes outside their training domain, remains a critical bottleneck in chemical discovery. This challenge is particularly acute in industrial discovery, where navigating uncharted chemical space to generate new intellectual property is a primary objective. Normal boiling points serve as a key benchmark for testing the extrapolative power of ML algorithms. A major limitation is that group-contribution methods are by design unable to generate predictions for molecules containing unparameterized fragments. Here, we demonstrate that this limitation can be overcome by replacing structural descriptors with thermodynamic properties computed directly from molecular dynamics simulations. We introduce a physics-augmented framework where a CatBoost regression model learns directly from ensemble-averaged cohesive energies, heats of vaporization, and densities extracted from atomistic liquid-phase simulations. Benchmark comparisons reveal that while both our physics-augmented model and conventional structure-based models perform comparably well on standard organic compounds, only the former maintains controlled error growth when extrapolating to structurally dissimilar chemical space. Our model successfully predicts boiling points for chemical classes entirely absent from training—including inorganic compounds, salts, and molecules with elements like Si, B, and Te—where structure-based models are fundamentally inapplicable. By encoding the intermolecular forces governing phase behavior, our framework establishes a generalizable strategy for property prediction beyond the structural boundaries of the existing methods.
@@ I Introduction
A rich spectrum of methods has been developed to address nBP and vapor-pressure prediction. Classical approaches, such as corresponding-states methods [63, 7], exploit reduced-property correlations based on critical data, but their scope is limited when such data is unavailable, or the target molecule’s chemotype diverges significantly [83]. Group-contribution (GC) schemes [27, 103, 20, 70, 8, 79, 74, 18, 89, 107], like the well-known Joback [49] or Nannoolal [75] methods, decompose molecules into predefined structural fragments. While rapid, and highly effective for molecules that can be segmented into their parameterized groups, GC methods are structurally blind to compounds built from elements or motifs not encoded in their parameterized fragment libraries [5]. In a similar vein, equation-of-state (EoS) models like Peng–Robinson or SAFT [81, 19, 38] can offer physically motivated predictions, but only when reliable component-specific parameters are known or sufficient experimental data is available for their parameterization. At the other end of the spectrum are first-principles methods, with COSMO-RS being a prominent example [56]. This approach uses quantum mechanics to predict thermodynamic properties, such as vapor pressure, for a wide range of chemicals—including organics, mixtures and ionic liquids—without requiring compound-specific experimental data [118, 98, 25]. However, as a continuum-based model, COSMO-RS averages intermolecular interactions, which leads to documented quantitative failures and large deviations for systems governed by strong, specific forces, such as those in ionic liquids [60, 77].
@@
The limitations of purely structural models motivate a direct solution: re-introducing the missing physics by using physics-based descriptors. This approach moves beyond the pre-defined structural descriptors—ranging from 2D fragments to 3D conformational properties—common in traditional QSPR, and instead computes features derived directly from first-principles simulations. One category of these descriptors is derived from quantum mechanics (QM), which provides a highly accurate description of a single molecule’s electronic structure [39, 13, 64]. However, while physically rigorous, these descriptors inherently describe a molecule in isolation, and thus neglect the crucial intermolecular interactions and ensemble effects that govern a condensed-phase property like the boiling point. A more direct approach is to use molecular dynamics (MD) simulations to explicitly model the molecular ensemble [35, 58, 97]. Indeed, computational studies have long leveraged MD to investigate phase transitions and related thermodynamic properties, from calculating vaporization enthalpies of complex fluids like ionic liquids [116] to characterizing liquid-vapor interfaces [41], and determining phase equilibria from rigorous but computationally demanding free energy calculations [68, 29]. However, these computational efforts have traditionally focused on either highly demanding, direct coexistence simulations to compute the nBP in the liquid-vapor equilibrium from scratch [78], or on using simulation data to develop more accurate force fields [76]. The computational expense of such simulations, however, has remained a significant barrier to their routine use for generating descriptors at scale, creating a need for more efficient methodologies. The challenge, therefore, lies in creating a hybrid framework that harnesses the predictive efficiency of data-driven models while being firmly anchored in the physical principles that govern molecular behavior. This goal of anchoring data-driven models in physical principles is central to the emerging paradigm of Physics-Informed Machine Learning (PIML) [50, 72], has gained traction across diverse scientific domains [51, 87, 119, 99, 111, 48].
@@ II.1 Training and Test Sets
The initial training dataset was sourced from the curated compilation in Ref.[54], which provides boiling point data for approximately 1,748 organic compounds (primarily hydrocarbons, alcohols, and amines) originally extracted from the CAS database [16]. To ensure the highest data quality, all experimental values were subsequently cross-verified against multiple authoritative databases: the NIST Chemistry WebBook[66], PubChem [55], and ChemSpider [80]. Compounds were excluded from our final dataset if they exhibited discrepancies greater than 10 K across sources, lacked corroborating entries, or were flagged with uncertainty markers. To ensure the physical validity of the simulations, a molecular weight threshold of 225 g/mol was applied. This step was necessary to exclude both low-molecular-weight compounds that would readily vaporize at the simulation temperatures (e.g., 300 K or 400 K) and high-molecular-weight compounds that would likely be solid or crystalline. In either case, the simulation would not represent a stable liquid-phase ensemble, invalidating the basis for our descriptor calculation. This filtering removed approximately 500 compounds, yielding a final training set of 1,280 entries with high-confidence experimental values. This collection forms the basis of our Training Set (available in section VI).
To rigorously evaluate model performance, two distinct test sets were curated, each designed to probe a different aspect of predictive capability. The first, a benchmark set for extrapolation, comprises active pharmaceutical ingredients (APIs) sourced from DrugBank [109]. These molecules were selected for their high structural complexity, often featuring more intricate scaffolds and a greater diversity of functional groups compared to the simpler, smaller molecules constituting the training set. These molecules, while structurally complex and diverse, are composed of standard organic elements. This composition ensures they fall within the theoretical applicability domain of modern chemoinformatics-based predictors, enabling a direct, head-to-head comparison of our model’s extrapolative performance against state-of-the-art methods on challenging, real-world structures.
The second, an out-of-domain challenge set, was curated to rigorously test the model’s generalization capabilities on chemical classes that lie far outside the applicability domain of traditional QSPR. It comprises chemotypes—including ionic liquids, salts, and molecules containing non-standard elements such as B, Si, and Te—for which many standard models are fundamentally inapplicable as they cannot generate a prediction for structures containing unparameterized elements or fragments. As a direct benchmark is therefore often impossible, predictions for this set are reported in order to characterize the model’s behavior and define the boundaries of its applicability in these challenging chemical spaces.
@@ II.2 Molecular Dynamics Simulations
For each compound and for each of the two force fields, three independent NPT simulations of 20 ns were performed at 300, 400 and 500 K, all at 1 atm with a 2 fs integration time step. Simulation boxes contained approximately 15,000 atoms. Thermodynamic properties were averaged over the final 5 ns of each trajectory. From these trajectories, a set of thermodynamic descriptors was computed, including the cohesive energy ($E_{\text{coh}}$), defined as the negative of the average intermolecular potential energy ($E_{\text{coh}}=-\langle U_{\text{inter}}\rangle$), heat of vaporization ($\Delta H_{\text{vap}}\approx E_{\text{coh}}+RT$), density ($\rho$), the Hildebrand solubility parameter ($\delta=\sqrt{(\Delta H_{\text{vap}}-RT)/V_{m}}$), and the isobaric specific heat capacity ($C_{P}$). The latter was computed from enthalpy fluctuations via the relation $C_{P}=(\langle H^{2}\rangle-\langle H\rangle^{2})/(k_{B}T^{2})$ [6].
@@ III Results
The boiling point of a substance is a direct manifestation of its intermolecular cohesive forces, a foundational principle of classical thermodynamics [9]. Based on this principle, we hypothesized that descriptors derived directly from the physics of the liquid phase could serve as a more robust foundation for property prediction than purely structural representations. Our approach therefore leverages cohesive energy, extracted directly from atomistic simulations, to encode the intermolecular interaction landscape that governs phase transitions, thereby avoiding reliance on predefined structural fragments or empirical group contributions.
@@
To validate this hypothesis, our initial objective was to establish a direct, quantitative relationship between the computed cohesive energy and experimental boiling points. We perform atomistic molecular dynamics simulations of 1,280 organic compounds (see section II.B) at three temperatures (300, 400, and 500K). To interrogate the robustness of this physics-based descriptor, we compare the performance of two structurally distinct force fields: OpenFF-2.0.0 [106], a modern open-source parameterization, and OPLS4 [67], an extensively validated commercial force field, yielding a total of 7,680 independent trajectories (1,280 compounds $\times$ 3 temperatures $\times$ 2 force fields). As shown in Fig.2, our analysis reveals a strong linear correlation between the simulated cohesive energy and experimental boiling points, though its quality is force-field dependent. This linear trend is physically grounded in Trouton’s rule, which implies a direct proportionality between the heat of vaporization ($\Delta H_{\text{vap}}$)—for which our cohesive energy is a proxy—and the boiling temperature ($T_{b}$[9]. Under the OpenFF-2.0.0 force field (Fig.2a), linear regression yields coefficients of determination ($R^{2}$) of 0.73 (300 K; blue), 0.76 (400 K; green), and 0.73 (500 K; pink). This correlation is further enhanced with the OPLS4 force field (Fig.2b), which affords $R^{2}$ values of 0.77 (300 K), 0.81 (400 K), and 0.82 (500 K). More important than the performance difference is the fact that a strong correlation exists for both, especially considering that neither force field was parameterized using boiling points from this dataset. This demonstrates the robustness of cohesive energy as a physics-based descriptor; its utility is not an artifact of a single, specific parameterization but stems from the underlying physical principle it represents. Taken together, this extensive linear analysis confirms that our physics-based approach robustly captures the fundamental intermolecular interactions governing boiling behavior. However, the residual scatter indicates that a simple linear model is insufficient. This is physically expected, as a linear model implicitly assumes a constant entropy of vaporization ($\Delta S_{\text{vap}}$) across all compounds (the basis of Trouton’s rule). The observed deviation from linearity thus arises from molecule-specific variations in this entropic contribution. This is precisely where a flexible, non-linear machine-learning model becomes essential, as it can learn the complex mapping from the enthalpic features to the final boiling point while implicitly accounting for these entropic variations.
@@
To contextualize the performance of our MD-only model, we conducted a systematic benchmark against the other architectures, presented in Figure 3B. The analysis reveals a clear performance hierarchy. Using the high-fidelity OPLS4 force field, the ”MD+Chemoinf.” hybrid model, which combines both MD and chemoinformatics features, attains the highest accuracy with a mean absolute error (MAE) of 6.2K and a root mean squared error (RMSE) of 11.7K (dark green bar). It is closely followed by the ”chemoinformatics-only” model, trained on a high-dimensional feature set of over 2,000 descriptors, which achieves an MAE of 6.9K and an RMSE of 12.2K (teal bar). Strikingly, our MD-only model (dark blue bar) remains highly competitive with an MAE of 8.2K and an RMSE of 13.8K. This result is highly significant. By replacing thousands of abstract structural features with a handful of physically meaningful descriptors, we reduce the feature dimensionality by over two orders of magnitude. This dramatic reduction is a standard strategy to mitigating the risk of overfitting, as the model is encouraged to learn the governing physical principles rather than memorizing spurious correlations from the training data [24, 10]. This, in turn, is fundamental to its enhanced extrapolative power. Remarkably, this robust generalization is achieved with only a minor sacrifice of 2K in MAE relative to the best-performing, high-dimensional model. The force field dependence is also instructive. The performance of the MD-only model is highly sensitive to the physical accuracy of the simulation, with the MAE degrading from 8.2K with OPLS4 to 12.2K with OpenFF-2.0.0 (hatched dark blue bar). In contrast, the hybrid model shows remarkable resilience, with its MAE only slightly increasing from 6.2K (OPLS4) to 6.9K (OpenFF). This suggests that structural descriptors can partially compensate for inaccuracies in simulation-derived properties, providing a practical pathway for deployment. For a final point of comparison, we include results from the Random Forest model reported by Kim et al. [54], the study from which our initial, uncurated dataset was sourced (light green bar) (see section S.II.A for more details about the data). This baseline model exhibits substantially higher errors (MAE = 19.9K, RMSE = 43.3K) than even our worst-performing variant. Several factors likely contribute to this performance gap. We hypothesize that two primary factors are: (i) our rigorous data curation produced a cleaner, more consistent training set, and (ii) our physics-based descriptors provide more causally relevant information than purely structural features. However, it must be acknowledged that a third confounding variable is the change in the learning algorithm itself, from a Random Forest to a CatBoost model, which may also contribute to the observed improvement. A full deconvolution of these effects is beyond the scope of this study.
@@
A deep dive into the feature importance, visualized in Figure 3C, reveals the distinct predictive logic of each model and highlights the efficiency of the physics-informed approach. Here, MD-derived features are colored in blue and chemoinformatics descriptors in red. For the MD-only model (top bar), we initially provided a pool of 15 physics-based features: five thermodynamic properties—cohesive energy ($E_{\text{coh}}$), heat of vaporization ($\Delta H_{\text{vap}}$), density ($\rho$), solubility parameter ($\delta$), and isobaric specific heat capacity ($C_{P}$)—each calculated at three different temperatures (see section S.II.B for further details). From this set, CatBoost’s internal feature selection algorithm identifies the most predictive, non-redundant subset. For the model trained on the OPLS4 data, we observe a highly compact feature selection, with the model identifying just three features as sufficient for optimal prediction. As shown in the top bar, the predictive power is overwhelmingly concentrated in the heat of vaporization at 300K ($\Delta H_{\text{vap}}$300K), which accounts for 84% of the model’s cumulative feature importance. The solubility parameter ($\delta$) at 300K (Solubility300K, 14%) and specific heat capacity at 500K (SpHeatCap500K, 2%) play minor, supporting roles. This specific three-feature selection, however, highlights the model’s data-driven nature rather than a fixed physical rule. The algorithm’s goal is to find the most predictive signal, and the composition of that signal is force-field dependent. For instance, when the same modeling process is applied to the features derived from the OpenFF simulations (see Supplementary Information, Fig.S1 in Supplementary Information), the model selects a larger and more varied set of six features. This is instructive, suggesting that for the OpenFF-derived data, the predictive signal is more distributed across several thermodynamic properties, prompting the algorithm to construct a more complex model to capture it effectively. What remains consistent is the flexibility of our strategy: the model successfully mines the available thermodynamic data, adapting its feature selection to the specific physical representation provided by each force field.
@@
In contrast, the chemoinformatics-based models leverage over 1000 features, but their importance distribution is revealing. In the chemoinformatics-only model, importance is led by molecular weight (52.1%), but a substantial portion (27.6%) is fragmented across a vast ”long tail” of abstract structural descriptors. This stark contrast in the number of employed features—a handful of physical properties versus over a thousand abstract ones—is a central finding of our work. The hybrid model, however, reveals the decisive influence of the physics-augmented features. In the OPLS4-based model, a single thermodynamic descriptor—the heat of vaporization at 300K ($\Delta H_{\text{vap}}$)—emerges as the overwhelmingly dominant feature, accounting for 61.7% of the total importance. This single physical property is nearly three times as important as the next most significant feature, molecular weight (20.7%), proving it provides the core predictive signal, which is then refined by structural information. This dominance of MD-derived features as the most significant predictive block is consistent across both force fields (see Supplementary Information, Fig.S1).

### 18 Mar 2026 v2

@@ Abstract
The limited extrapolative power of structure-based machine learning (ML) models is a critical bottleneck in chemical discovery, particularly for industrial R&D, where navigating uncharted chemical space to find next-generation materials or drugs is paramount. These models, reliant on structural descriptors or graph neural networks (GNNs), often fail when predicting properties for molecules with novel chemotypes. Here, we introduce a physics-augmented ML framework that overcomes this limitation. Our approach replaces conventional structural inputs with thermodynamic properties such as cohesive energy, heat of vaporization, and density, derived directly from molecular dynamics (MD) simulations. While performing comparably to structure-based models on known organic compounds, our method uniquely maintains low error when extrapolating to dissimilar chemical spaces. Crucially, it accurately predicts boiling points for entire chemical classes absent from the training set, including inorganic compounds, salts, and molecules with elements such as Si, B, and Te. By learning from the intermolecular forces that govern phase transitions, our framework provides a more fundamental and generalizable strategy for molecular property prediction, enabling chemical exploration beyond established structural domains.
@@ I Introduction
A rich spectrum of methods has been developed to address nBP and vapor-pressure prediction. Classical approaches, such as corresponding-states methods [64, 7], exploit reduced-property correlations based on critical data, but their scope is limited when such data is unavailable, or the target molecule’s chemotype diverges significantly [84]. Group-contribution (GC) schemes [27, 105, 20, 71, 8, 80, 75, 18, 90, 109], like the well-known Joback [50] or Nannoolal [76] methods, decompose molecules into predefined structural fragments. While rapid, and highly effective for molecules that can be segmented into their parameterized groups, GC methods are structurally blind to compounds built from elements or motifs not encoded in their parameterized fragment libraries [5]. In a similar vein, equation-of-state (EoS) models like Peng–Robinson or SAFT [82, 19, 38] can offer physically motivated predictions, but only when reliable component-specific parameters are known or sufficient experimental data is available for their parameterization.
@@
The limitations of purely structural models motivate a direct solution: re-introducing the missing physics by using physics-based descriptors. This approach moves beyond the pre-defined structural descriptors—ranging from 2D fragments to 3D conformational properties—common in traditional QSPR, and instead computes features derived directly from first-principles simulations. One category of these descriptors is derived from quantum mechanics (QM), which provides a highly accurate description of a single molecule’s electronic structure [39, 13, 65]. However, while physically rigorous, these descriptors inherently describe a molecule in isolation, and thus neglect the crucial intermolecular interactions and ensemble effects that govern a condensed-phase property like the boiling point, a subtlety that can lead to misleading conclusions if not interpreted carefully [43]. A more sophisticated approach, COSMO-RS, attempts to bridge this gap [57]. It uses QM to predict thermodynamic properties for a wide range of chemicals by placing the molecule in a virtual conductor, thereby modeling the ensemble effect via a continuum [120, 99, 25]. Yet, as a continuum-based model, COSMO-RS averages intermolecular interactions, leading to documented quantitative failures for systems governed by strong, specific forces, such as those in ionic liquids [61, 78]. A more direct approach is to use molecular dynamics (MD) simulations to explicitly model the molecular ensemble [35, 59, 98]. Indeed, computational studies have long leveraged MD to investigate phase transitions and related thermodynamic properties, from calculating vaporization enthalpies of complex fluids like ionic liquids [118] to characterizing liquid-vapor interfaces [41], and determining phase equilibria from rigorous but computationally demanding free energy calculations [69, 29]. However, these computational efforts have traditionally focused on either highly demanding, direct coexistence simulations to compute the nBP in the liquid-vapor equilibrium from scratch [79], or on using simulation data to develop more accurate force fields [77]. The computational expense of such simulations, however, has remained a significant barrier to their routine use for generating descriptors at scale, creating a need for more efficient methodologies. The challenge, therefore, is to create a hybrid framework that harnesses the predictive efficiency of data-driven models while being firmly anchored in the physical principles that govern molecular behavior. This goal of anchoring data-driven models in physical principles is central to the emerging paradigm of Physics-Informed Machine Learning (PIML) [51, 73, 52, 88, 121, 100, 113, 49].
@@ II.1 Training and Test Sets
The initial training dataset was sourced from the curated compilation in Ref.[55], which provides boiling point data for approximately 1,748 organic compounds (primarily hydrocarbons, alcohols, and amines) originally extracted from the CAS database [16]. To ensure the highest data quality, all experimental values were subsequently cross-verified against multiple authoritative databases: the NIST Chemistry WebBook[67], PubChem [56], and ChemSpider [81]. Compounds were excluded from our final dataset if they exhibited discrepancies greater than 10 K across sources, lacked corroborating entries, or were flagged with uncertainty markers. To ensure the physical validity of the simulations, a molecular weight threshold of 225 g/mol was applied. This step was necessary to exclude high-molecular-weight compounds that would likely be solid or crystalline. In either case, the simulation would not represent a stable liquid-phase ensemble, invalidating the basis for our descriptor calculation. This filtering removed approximately 500 compounds, yielding a final training set of 1,280 entries with high-confidence experimental values. This collection forms the basis of our Training Set (available in section VI).
To rigorously evaluate model performance, two distinct test sets were curated, each designed to probe a different aspect of predictive capability. The first, a benchmark set for extrapolation, comprises active pharmaceutical ingredients (APIs) sourced from DrugBank [111]. These molecules were selected for their high structural complexity, often featuring more intricate scaffolds and a greater diversity of functional groups compared to the simpler, smaller molecules constituting the training set. These molecules, while structurally complex and diverse, are composed of standard organic elements. This composition ensures they fall within the theoretical applicability domain of modern GNNs, enabling a direct, head-to-head comparison of our model’s extrapolative performance against state-of-the-art methods on challenging, real-world structures.
The second, an out-of-domain challenge set, was curated to challenge the model’s generalization capabilities on chemical classes that lie far outside the applicability domain of traditional QSPR. It comprises chemotypes—including ionic liquids, salts, and molecules containing non-standard elements such as B, Si, and Te—for which many standard models are fundamentally inapplicable as they cannot generate a prediction for structures containing unparameterized elements or fragments. As a direct benchmark is therefore often impossible, predictions for this set are reported in order to characterize the model’s behavior and define the boundaries of its applicability in these challenging chemical spaces.
@@ II.2 Molecular Dynamics Simulations
For each compound and for each of the two force fields, three independent NPT simulations of 20 ns were performed at 300, 400 and 500 K, all at 1 atm with a 2 fs integration time step. Simulation boxes contained approximately 15,000 atoms. Thermodynamic properties were averaged over the final 5 ns of each trajectory. From these trajectories, a set of thermodynamic descriptors was computed, including the cohesive energy ($E_{\text{coh}}$), defined as the negative of the average intermolecular potential energy ($E_{\text{coh}}\approx-\langle U_{\text{inter}}\rangle$), heat of vaporization ($\Delta H_{\text{vap}}\approx E_{\text{coh}}+RT$), density ($\rho$), the Hildebrand solubility parameter ($\delta=\sqrt{(\Delta H_{\text{vap}}-RT)/V_{m}}$), and the isobaric specific heat capacity ($C_{P}$). The latter was computed from enthalpy fluctuations via the relation $C_{P}=(\langle H^{2}\rangle-\langle H\rangle^{2})/(k_{B}T^{2})$ [6].
@@ III Results
The boiling point of a substance is a direct manifestation of its intermolecular cohesive forces, a foundational principle of classical thermodynamics [9]. Based on this principle, we hypothesized that descriptors derived directly from the physics of the liquid phase could serve as a more robust foundation for property prediction than purely structural representations. Our approach therefore leverages cohesive energy, extracted directly from atomistic simulations, to encode the intermolecular interaction landscape that governs phase transitions.
@@
To validate this hypothesis, our initial objective was to establish a direct, quantitative relationship between the computed cohesive energy and experimental boiling points. We perform atomistic molecular dynamics simulations of 1,280 organic compounds (see section II.B) at three temperatures (300, 400, and 500K). To interrogate the robustness of this physics-based descriptor, we compare the performance of two structurally distinct force fields: OpenFF-2.0.0 [108], a modern open-source parameterization, and OPLS4 [68], an extensively validated commercial force field, yielding a total of 7,680 independent trajectories (1,280 compounds $\times$ 3 temperatures $\times$ 2 force fields). As shown in Fig.2, our analysis reveals a strong linear correlation between the simulated cohesive energy and experimental boiling points, though its quality is force-field dependent [102]. This linear trend is physically grounded in Trouton’s rule, which implies a direct proportionality between the heat of vaporization ($\Delta H_{\text{vap}}$)—for which our cohesive energy is a proxy—and the boiling temperature ($T_{b}$[9]. Under the OpenFF-2.0.0 force field (Fig.2a), linear regression yields coefficients of determination ($R^{2}$) of 0.73 (300 K; blue), 0.76 (400 K; green), and 0.73 (500 K; pink). This correlation is further enhanced with the OPLS4 force field (Fig.2b), which affords $R^{2}$ values of 0.77 (300 K), 0.81 (400 K), and 0.82 (500 K). More important than the performance difference is the fact that a strong correlation exists for both, especially considering that neither force field was parameterized using boiling points from this dataset. This demonstrates the robustness of cohesive energy as a physics-based descriptor; its utility is not an artifact of a single, specific parameterization but stems from the underlying physical principle it represents. Taken together, this extensive linear analysis confirms that our physics-based approach robustly captures the fundamental intermolecular interactions governing boiling behavior. However, the residual scatter indicates that a simple linear model is insufficient. This is physically expected, as a linear model implicitly assumes a constant entropy of vaporization ($\Delta S_{\text{vap}}$) across all compounds (the basis of Trouton’s rule). The observed deviation from linearity thus arises from molecule-specific variations in this entropic contribution. This is precisely where a flexible, non-linear machine-learning model becomes essential, as it can learn the complex mapping from the enthalpic features to the final boiling point while implicitly accounting for these entropic variations.
@@
To contextualize the performance of our MD-only model, we conducted a systematic benchmark against the other architectures, presented in Figure 3B. The analysis reveals a clear performance hierarchy. Using the high-fidelity OPLS4 force field, the ”MD+Chemoinf.” hybrid model, which combines both MD and chemoinformatics features, attains the highest accuracy with a mean absolute error (MAE) of 6.2K and a root mean squared error (RMSE) of 11.7K (dark green bar). It is closely followed by the ”chemoinformatics-only” model, trained on a high-dimensional feature set of over 2,000 descriptors, which achieves an MAE of 6.9K and an RMSE of 12.2K (teal bar). Strikingly, our MD-only model (dark blue bar) remains highly competitive with an MAE of 8.2K and an RMSE of 13.8K. This result is highly significant. By replacing thousands of abstract structural features with a handful of physically meaningful descriptors, we reduce the feature dimensionality by over two orders of magnitude. This dramatic reduction is a standard strategy to mitigating the risk of overfitting, as the model is encouraged to learn the governing physical principles rather than memorizing spurious correlations from the training data [24, 10]. This, in turn, is fundamental to its enhanced extrapolative power. Remarkably, this robust generalization is achieved with only a minor sacrifice of 2K in MAE relative to the best-performing, high-dimensional model. The force field dependence is also instructive. The performance of the MD-only model is highly sensitive to the physical accuracy of the simulation, with the MAE degrading from 8.2K with OPLS4 to 12.2K with OpenFF-2.0.0 (hatched dark blue bar). In contrast, the hybrid model shows remarkable resilience, with its MAE only slightly increasing from 6.2K (OPLS4) to 6.9K (OpenFF). This suggests that structural descriptors can partially compensate for inaccuracies in simulation-derived properties, providing a practical pathway for deployment. For a final point of comparison, we include results from the Random Forest model reported by Kim et al. [55], the study from which our initial, uncurated dataset was sourced (light green bar) (see section S.II.A for more details about the data). This baseline model exhibits substantially higher errors (MAE = 19.9K, RMSE = 43.3K) than even our worst-performing variant. Several factors likely contribute to this performance gap. We hypothesize that two primary factors are: (i) our rigorous data curation produced a cleaner, more consistent training set, and (ii) our physics-based descriptors provide more causally relevant information than purely structural features. However, it must be acknowledged that the observed improvement cannot be attributed solely to the descriptors. The overall modeling strategy itself introduces confounding variables. These include not only the change in the learning algorithm (from Random Forest to CatBoost), but also the distinct feature selection and hyperparameter optimization processes applied to each model. A full deconvolution of these effects is beyond the scope of this study.
@@
A deep dive into the feature importance, visualized in Figure 3C, reveals the distinct predictive logic of each model and highlights the efficiency of the physics-informed approach. Here, MD-derived features are colored in blue and chemoinformatics descriptors in red. For the MD-only model (top bar), we initially provided a pool of 15 physics-based features: five thermodynamic properties—cohesive energy ($E_{\text{coh}}$), heat of vaporization ($\Delta H_{\text{vap}}$), density ($\rho$), solubility parameter ($\delta$), and isobaric specific heat capacity ($C_{P}$)—each calculated at three different temperatures (see section S.II.B for further details). From this set, CatBoost’s internal feature selection algorithm identifies the most predictive, non-redundant subset. For the model trained on the OPLS4 data, we observe a highly compact feature selection, with the model identifying just three features as sufficient for optimal prediction. As shown in the top bar, the predictive power is overwhelmingly concentrated in the heat of vaporization at 300K ($\Delta H_{\text{vap}}$300K), which accounts for 84% of the model’s cumulative feature importance. The solubility parameter ($\delta$) at 300K (Solubility300K, 14%) and specific heat capacity at 500K (SpHeatCap500K, 2%) play minor, supporting roles. This specific three-feature selection, however, highlights the model’s data-driven nature rather than a fixed physical rule. The algorithm’s goal is to find the most predictive signal, and the composition of that signal is force-field dependent. For instance, when the same modeling process is applied to the features derived from the OpenFF simulations (see Supplementary Information, Fig.S1 in Supplementary Information), the model selects a larger and more varied set of six features. This is instructive, suggesting that for the OpenFF-derived data, the predictive signal is more distributed across several thermodynamic properties, prompting the algorithm to construct a more complex model to capture it effectively.
@@
In contrast, the chemoinformatics-based models leverage over 1000 features, but their importance distribution is revealing. In the chemoinformatics-only model, importance is led by molecular weight (52.1%), but a substantial portion (27.6%) is fragmented across a vast ”long tail” of abstract structural descriptors. A stark contrast emerges in the number of employed features: our model relies on a handful of physical properties, whereas the baseline requires over a thousand abstract ones. The hybrid model, however, reveals the decisive influence of the physics-augmented features. In the OPLS4-based model, a single thermodynamic descriptor—the heat of vaporization at 300K ($\Delta H_{\text{vap}}$)—emerges as the overwhelmingly dominant feature, accounting for 61.7% of the total importance. This single physical property is nearly three times as important as the next most significant feature, molecular weight (20.7%), proving it provides the core predictive signal, which is then refined by structural information. This dominance of MD-derived features as the most significant predictive block is consistent across both force fields (see Supplementary Information, Fig.S1).